A +100μC charged is place at the origin. A +200μC charge is place 2m above the origin and a -300μC charged is placed on 3m to the right of the origin. What is the magnitude and direction of the net force acting on the charge placed at the origin?
1
Expert's answer
2021-08-04T04:40:37-0400


The charges and forces are shown in the figure. Let "q_1 = 100\mu C = 10^{-4}C", "q_2 = 2\times 10^{-4} C", "q_3 = 3\times 10^{-4} C".

Then the forces according to the Coulom's law are:


"F_{12} = \dfrac{9\times 10^9\cdot 10^{-4}\cdot 2\times 10^{-4}}{2^2} = 45N"

"F_{13} = \dfrac{9\times 10^9\cdot 10^{-4}\cdot 3\times 10^{-4}}{3^2} = 3N"

Since the angle betwen them is right one, the magnitude of the net force can be found from Pyphagorean's theorem:


"F = \sqrt{45^2 + 3^2} \approx 45.1N"

The direction is:


"\theta = 360\degree - \arctan\left( \dfrac{F_{12}}{F_{13}} \right)=360\degree - \arctan\left( \dfrac{45}{3} \right) \approx 274\degree"

Answer. 45.1 N, 274 degrees with positive x-axis.